A frictionless free piston-cylinder device contains 5125 g of CH4 (methane) at -70.0°C. The initial volume the CH4 occupies iQuestion 2 (10 Marks) A frictionless free piston-cylinder device contains 8.60 kg of water at 450 kPa. The initial volume the1a) 701 K 1b) 557 K 1c) Key points to have in answer: i) Relating Z not being to close unity, ii) ideal gas large error, iii)please show work
Answer #1

Solution to question (1).

Given m=5125 g T,= - 70 c= 10+273-15=203.15 = 145x10 m - soox 103 m3 T2 = ? (a) By using ideal gas equation. PV = P₂ V₂ MRT M

Reduced pressure: - PR= R - 3072 ² = 0.8019 Per 4.64 203-15 19. - 1.063 Teto 203.15 1.063 using charts for methane, at PR = 0